Littlefoot a 2 kg DSH needs the pressor dopamine to treat her hypotension. The dose is 5 mcg/kg/min to start. The concentration of dopamine is 40 mg/ml. The rate is 5 ml/hr. How many mLs of dopamine will Littlefoot need if the CRI is made up for 24 hours ? What is the total volume to be made
Answer #1

Dopamine - 5 mcg \times 2 = 10 mcg/ min

For 1 hour 10 mcg \times 60 = 600 mcg

For 24 hours = 600 mcg \times 24 = 14400 mcg for a day

14400 mcg = 14.4 mg

Required medication = required dose \times available volume \div available dose

                                  = 14.4 \times 1 \div 40

                                  = 0.36 ml of dopamine or 0.4 ml of dopamine

The rate is 5 ml per hour , for 24 hours = 5 \times 24 = 120 ml

120 ml - 0.36 ml = 119.64 ml of diluent or 119.6 ml of diluent